Solve by completing the square
x^2 = 5x + 2

Thanks

Answer by Nate(3500)

x^2 = 5x + 2
x^2 - 5x = 2
(x - 5/2)^2 = 8/4 + 25/4
(x - 5/2)^2 = 33/4
x - 5/2 = +-sqrt(33)/2
x = (5 +- sqrt(33))/2